A carnot engine takes 3*10to the6 cal of heat from a reservoir at 6*27C, and gives it to a sink at 27
C. The work done by the engine is

Grade:12

2 Answers

Saurabh Koranglekar
3 years ago
Dear student

Please attach the image of the question

Regards
Arun
3 years ago

T1=273+627=900k

T2=273+27=300

N=W/Q1=1-T2/T1

W/3000k cal=1–300/900

W=2000kcal

2000×4.2KJ=8.4×10^6 joule